write the dimensions of a nd b in the relation P=b-x^2/at when P=power, x=distance,, t=time ?

Grade:12th pass

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The dimension of P is ML^{2}T^{-3}.
Hence the dimension of b is ML^{2}T^{-3}.
The dimension of a is thus M^{-1}T^{2}.
Regards
Sumit
Abhishek Nerkar
6 years ago
P=b-x^2/at.....1As x^2 is subtract from b, therefore the dimensions of b=(L^2)We can write relation 1 asP=(L^2)/at=(L^2)/(L^2M^1T^-3)×(T^2)Therefore a=(L^0M^1T^2)This,a=(L^0M^1T^2) and b=(L^2)